### v4

@@ 3.1 Discounted Infinite-horizon MDP
We consider a discounted infinite-horizon MDP, denoted by a quintuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}=\{1,2,...,|\mathcal{S}|\}$ and $\mathcal{A}=\{1,2,...,|\mathcal{A}|\}$ represent the finite state space and action space, respectively. The probability transition kernel $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[-1,1]$ determines the likelihood $P\left(s^{\prime}\mid s,a\right)$ of transitioning from state $s$ to state $s^{\prime}$ upon taking action $a$. The reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ assigns an immediate reward $r(s,a)$ for executing action $a$ in state $s$. Additionally, the discount factor $\gamma\in(0,1)$ adjusts the weighting of future rewards.
@@ 3.2 Policy, and $Q$-function
$Q^{\pi}(s,a)=r(s,a)+\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r\left(s_{k},a_{k}\right)\mid s_{0}=s,a_{0}=a\right],\forall\leavevmode\nobreak\ (s,a)\in\mathcal{S}\times\mathcal{A}.$ (2)
@@ 5.1 CompFedRL with UnbiasedComp
where $\rho=1-\beta+\beta(1-\eta)^{K}$, $e_{1}=\frac{4\gamma}{C}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\right)$, $e_{2}=\frac{1}{(1-(1-\eta)^{K})}\left(\sqrt{16\frac{4q_{2}||\mathcal{S}|\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$, and $C=(1-\gamma)(1-(1-\eta)^{K})$.
@@
where $\bar{e}_{1}=\frac{4\gamma}{(1-\gamma)}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\right)$, and
$\bar{e}_{2}=\left(\sqrt{16\frac{4q_{2}||\mathcal{S}|\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$.
@@
From this bound, CompFedRL with $K=1$ and UnbiasedComp close to the identity operator (i.e. $q_{2}\rightarrow 0$ and $q_{\infty}\rightarrow 0$) and with $I\geq\max\left(B_{1},B_{2}\right)$ where $B_{1}=\left(\frac{4\gamma}{1-\gamma}\right)^{2}\frac{16\log(4|\mathcal{S}||\mathcal{A}|TK/\delta)}{\epsilon^{2}}$ and $B_{2}=\left(\frac{4\gamma}{1-\gamma}\right)\frac{4\sqrt{\log(4|\mathcal{S}||\mathcal{A}|TK/\delta)}}{\epsilon}$ achieves $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$ iterations, when $\gamma\leq\frac{\epsilon}{8+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$, then we have the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity that is close to the lower bound for a single-agent $Q$-learning by [39].
@@ 5.2 CompFedRL with BiasedComp
From Theorem 5.2 CompFedRL with BiasedComp converges linearly with high probability with three residual error terms. The following observations from this theorem are similar to those from Theorem 5.1 First, there is a trade-off between convergence speed and solution accuracy of the algorithm when we adjust the learning rate $\eta$, the federated parameter $\beta$, and the number of local epochs $K$. In particular, the algorithm with $K\gg 1$ and $\beta,\eta\rightarrow 1$ achieves high speed at the price of low solution accuracy. Second, due to the dependency of the first error term in the convergence bound on $1/\sqrt{I}$, CompFedRL with BiasedComp achieves the speedup with respect to the number of agents similarly to CompFedRL with UnbiasedComp in Theorem 5.1 and to multi-agent TD(0) learning using error-feedback analyzed by [47]. Third, the last error term is due to the compression $\alpha$ and can be diminished by decreasing $\beta$ and increasing $\eta,K$, whereas the second error term is negligible when the discount factor $\gamma$ is close to zero. Fourth, we obtain the convergence for full-precision synchronous $Q$-learning when we let $\alpha=1$ in Theorem 5.2.
@@
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta\eta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{\eta I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\gamma}{\eta(1-\gamma)}.$
@@
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}e_{1}+\frac{2\gamma}{1-\gamma}.$
In conclusion, this algorithm reaches $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$ iterations when $I=\frac{16}{(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}$ and $\gamma\leq\frac{\epsilon}{6+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$, then we achieve the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity which nearly matches the lower bound for the single-agent $Q$-learning by [39].
@@ 0.A.1 Useful lemmas
$\Delta_{t,k+1}=(1-\eta)\Delta_{t,k}+\frac{\eta}{I}\sum_{i=1}^{I}[r+\gamma P^{(i)}_{t,k+1}V^{(i)}_{t,k}],$
@@
where $E^{1}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}(P_{t,k+1}^{(i)}-P)V^{(i)}_{t}$ and $E^{2}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}P(V^{(i)}_{t,k}-V^{\star})$.
To complete the proof, we must bound $\|E_{1}\|_{\infty}$ and $\|E_{2}\|_{\infty}$. We first bound we bound $\|E_{2}\|_{\infty}$. By the fact that $\|P(s,a)\|_{1}\leq 1$, $\|V^{(i)}_{t}\|_{\infty}\leq\frac{1}{1-\gamma}$, and $\|V^{\star}\|_{\infty}$,
$\|E_{2}\|_{\infty}$$\leq$ $\frac{\gamma\eta}{I}\sum_{t=0}^{T-1}(1-\eta)^{T-1-l}\sum_{i=1}^{I}\|P\|_{1}(\|V^{(i)}_{t}\|_{\infty}+\|V^{\star}\|_{\infty})$ (10)
$\leq$ $\frac{2\gamma}{1-\gamma}\sum_{t=0}^{T-1}\eta(1-\eta)^{T-1-l}\leq\frac{2\gamma}{1-\gamma}.$
Next, we bound $\|E_{1}\|_{\infty}$ by Lemma 4 with $w_{t}^{(i)}=(P_{t+1}^{(i)}-P)V^{(i)}_{t}$. By the fact that $w_{t}^{(i)}$ is independent of the transition events of other agents, and that $\|P^{(i)}_{t}(s,a)\|_{1}\leq 1$, $\|P(s,a)\|_{1}\leq 1$, and $\|V^{(i)}_{t}\|_{\infty}\leq\frac{1}{1-\gamma}$, we can prove by proof arguments in Lemma 1 of [68] that $W=\frac{2}{1-\gamma}$ and $\sigma^{2}=\frac{2}{(1-\gamma)^{2}}$. Therefore, for every $s,a,t$
$\|E_{1}\|_{\infty}$$\leq$ $\gamma\left(\sqrt{\frac{16\eta}{I(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}+\frac{8}{3I(1-\gamma)}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}\right)$ (11)
@@
where $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$. Since $\Delta_{0}=\overline{Q}_{t}$, we complete the proof.
@@ 0.A.1.2 Proof of Lemma 2
$\sum_{i=1}^{I}\mathbf{E}(e_{t}^{(i)})^{2}$$\leq$ $\sum_{i=1}^{I}q_{2}\|Q^{(i)}_{t,K}-\overline{Q}_{t}\|_{2}^{2}\leq\frac{4Iq_{2}||\mathcal{S}|\mathcal{A}|}{(1-\gamma)^{2}}.$
@@
$\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}\leq\frac{1}{\sqrt{I}}\left(\sqrt{16\frac{4q_{2}||\mathcal{S}|\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right).$
@@
with probability at least $1-\delta/T$. Here, $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}$ and $e_{2}=\sqrt{16\frac{4q_{2}||\mathcal{S}|\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}$. Finally from Lemma 3, with probability at least $1-\delta$
@@
$\hat{\Delta}_{t+1}=(1-\beta)\hat{\Delta}_{t}+\beta\left(\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right)+\beta^{2}\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}.$

### v6

@@ 3.1 Discounted Infinite-horizon MDP
We consider a discounted infinite-horizon MDP, denoted by a quintuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}=\{1,2,...,|\mathcal{S}|\}$ and $\mathcal{A}=\{1,2,...,|\mathcal{A}|\}$ represent the finite state space and action space, respectively. The transition probability kernel $P:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ determines the likelihood $P\left(s^{\prime}\mid s,a\right)$ of transitioning from state $s$ to state $s^{\prime}$ upon taking action $a$. The reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow[-1,1]$ assigns an immediate reward $r(s,a)$ for executing action $a$ in state $s$. Additionally, the discount factor $\gamma\in(0,1)$ adjusts the weighting of future rewards.
@@ 3.2 Policy, and $Q$-function
$Q^{\pi}(s,a)=\mathbb{E}\left[\sum_{k=0}^{\infty}\gamma^{k}r\left(s_{k},a_{k}\right)\mid s_{0}=s,a_{0}=a\right],\forall\leavevmode\nobreak\ (s,a)\in\mathcal{S}\times\mathcal{A}.$ (2)
@@ 5.1 CompFedRL with UnbiasedComp
where $\rho=1-\beta+\beta(1-\eta)^{K}$, $e_{1}=\frac{4\gamma}{C}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}\right)$, $e_{2}=\frac{1}{(1-(1-\eta)^{K})}\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$, and $C=(1-\gamma)(1-(1-\eta)^{K})$.
@@
where $\bar{e}_{1}=\frac{4\gamma}{(1-\gamma)}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|T}{\delta}}\left(1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|T}{\delta}}\right)$, and
$\bar{e}_{2}=\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right)$.
@@
From this bound, CompFedRL with $K=1$ and UnbiasedComp close to the identity operator (i.e. $q_{2}\rightarrow 0$ and $q_{\infty}\rightarrow 0$) and with $I\geq\max\left(B_{1},B_{2}\right)$ where $B_{1}=\left(\frac{4\gamma}{1-\gamma}\right)^{2}\frac{16\log(4|\mathcal{S}||\mathcal{A}|T/\delta)}{\epsilon^{2}}$ and $B_{2}=\left(\frac{4\gamma}{1-\gamma}\right)\frac{4\sqrt{\log(4|\mathcal{S}||\mathcal{A}|T/\delta)}}{\epsilon}$ achieves $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$ iterations, when $\gamma\leq\frac{\epsilon}{8+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(4\|\overline{Q}_{0}-Q^{\star}\|_{\infty}/\epsilon)$, then we have the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity that is close to the lower bound for a single-agent $Q$-learning by [39].
@@ 5.2 CompFedRL with BiasedComp
From Theorem 5.2 CompFedRL with BiasedComp converges linearly with high probability with three residual error terms. The following observations from this theorem are similar to those from Theorem 5.1. First, there is a trade-off between convergence speed and solution accuracy of the algorithm when we adjust the learning rate $\eta$, the federated parameter $\beta$, and the number of local epochs $K$. In particular, the algorithm with $K\gg 1$ and $\beta,\eta\rightarrow 1$ achieves high speed at the price of low solution accuracy. Second, due to the dependency of the first error term in the convergence bound on $1/\sqrt{I}$, CompFedRL with BiasedComp achieves the speedup with respect to the number of agents similarly to CompFedRL with UnbiasedComp in Theorem 5.1 and to multi-agent TD(0) learning using error-feedback analyzed by [47]. Third, the last error term is due to the compression $\alpha$ and can be diminished by decreasing $\beta$ and increasing $\eta,K$, whereas the second error term is negligible when the discount factor $\gamma$ is close to zero. Fourth, we obtain the convergence for full-precision synchronous $Q$-learning when we let $\alpha=1$ in Theorem 5.2.
@@
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta\eta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{\eta I}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}e_{1}+\frac{2\gamma}{\eta(1-\gamma)}.$
@@
$\|\overline{Q}_{T}-Q^{\star}\|_{\infty}$$\leq$ $(1-\beta)^{T}\|\overline{Q}_{0}-Q^{\star}\|_{\infty}+\frac{4}{1-\gamma}\sqrt{\frac{1}{I}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}e_{1}+\frac{2\gamma}{1-\gamma}.$
In conclusion, this algorithm reaches $\|\overline{Q}_{T}-Q^{\star}\|_{\infty}\leq\epsilon$ within $T=\frac{1}{\beta}\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$ iterations when $I=\frac{16}{(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}$ and $\gamma\leq\frac{\epsilon}{6+\epsilon}$. If $\beta=(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})/\log(3\|\overline{Q}_{0}-Q^{\star}\|/\epsilon)$, then we achieve the $\mathcal{O}\left(\frac{1}{(1-\gamma)^{4}\min(\epsilon,\epsilon^{2})}\right)$ complexity which nearly matches the lower bound for the single-agent $Q$-learning by [39].
@@ 0.A.1 Useful lemmas
$\Delta_{t,k+1}=(1-\eta)\Delta_{t,k}+\frac{\eta}{I}\sum_{i=1}^{I}[r+\gamma P^{(i)}_{t,k+1}V^{(i)}_{t,k}-Q^{\star}],$
@@
where $E^{1}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}(P_{t,k+1}^{(i)}-P)V^{(i)}_{t,k}$ and $E^{2}_{t}=\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}P(V^{(i)}_{t,k}-V^{\star})$.
To complete the proof, we must bound $\|E^{1}_{t}\|_{\infty}$ and $\|E^{2}_{t}\|_{\infty}$. We first bound we bound $\|E^{2}_{t}\|_{\infty}$. By the fact that $\|P(s,a)\|_{1}\leq 1$, $\|V^{(i)}_{t,k}\|_{\infty}$, and $\|V^{\star}\|_{\infty}\leq\frac{1}{1-\gamma}$,
$\|E^{2}_{t}\|_{\infty}$$\leq$ $\frac{\gamma\eta}{I}\sum_{k=0}^{K-1}(1-\eta)^{K-1-k}\sum_{i=1}^{I}\|P\|_{1}(\|V^{(i)}_{t,k}\|_{\infty}+\|V^{\star}\|_{\infty})$ (10)
$\leq$ $\frac{2\gamma}{1-\gamma}\sum_{k=0}^{K-1}\eta(1-\eta)^{K-1-k}\leq\frac{2\gamma}{1-\gamma}.$
Next, we bound $\|E^{1}_{t}\|_{\infty}$ by Lemma 4 with $w_{k}^{(i)}=(P_{t,k+1}^{(i)}-P)V^{(i)}_{t,k}$. By the fact that $w_{k}^{(i)}$ is independent of the transition events of other agents, and that $\|P^{(i)}_{t,k}(s,a)\|_{1}\leq 1$, $\|P(s,a)\|_{1}\leq 1$, and $\|V^{(i)}_{t,k}\|_{\infty}\leq\frac{1}{1-\gamma}$, we can prove by proof arguments in Lemma 1 of [68] that $W=\frac{2}{1-\gamma}$ and $\sigma^{2}=\frac{2}{(1-\gamma)^{2}}$. Therefore, for every $s,a,t$
$\|E^{1}_{t}\|_{\infty}$$\leq$ $\gamma\left(\sqrt{\frac{16\eta}{I(1-\gamma)^{2}}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}}+\frac{8}{3I(1-\gamma)}\log\frac{2|\mathcal{S}||\mathcal{A}|T}{\delta}\right)$ (11)
@@
where $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{2|\mathcal{S}||\mathcal{A}|TK}{\delta}}$. Since $\Delta_{t,0}=\overline{Q}_{t}-Q^{\star}$, we complete the proof.
@@ 0.A.1.2 Proof of Lemma 2
$\sum_{i=1}^{I}\mathbf{E}(e_{t}^{(i)})^{2}$$\leq$ $\sum_{i=1}^{I}q_{2}\|Q^{(i)}_{t,K}-\overline{Q}_{t}\|_{2}^{2}\leq\frac{4Iq_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}.$
@@
$\left\|\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t}\right\|_{\infty}\leq\frac{1}{\sqrt{I}}\left(\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}\right).$
@@
with probability at least $1-\delta/T$. Here, $e_{1}=1+\frac{1}{\sqrt{\eta I}}\sqrt{\log\frac{4|\mathcal{S}||\mathcal{A}|TK}{\delta}}$ and $e_{2}=\sqrt{16\frac{4q_{2}|\mathcal{S}||\mathcal{A}|}{(1-\gamma)^{2}}\log\frac{4T}{\delta}}+\frac{4}{3}\frac{2q_{\infty}\sqrt{I}}{1-\gamma}\log\frac{4T}{\delta}$. Finally from Lemma 3, with probability at least $1-\delta$
@@
$\hat{\Delta}_{t+1}=(1-\beta)\hat{\Delta}_{t}+\beta\left(\frac{1}{I}\sum_{i=1}^{I}Q^{(i)}_{t,K}-Q^{\star}\right)+\beta^{2}\frac{1}{I}\sum_{i=1}^{I}e^{(i)}_{t},$
where $e^{(i)}_{t+1}=Q_{t,K}^{(i)}-\overline{Q}_{t}+e^{(i)}_{t}-\textsf{Compress}(Q_{t,K}^{(i)}-\overline{Q}_{t}+e^{(i)}_{t})$. From the definition of the $\ell_{\infty}$-norm and by the triangle inequality,